Solve the system by addition.
5x – 3y = 13
4x – 3y = 11

Answer by vertciel(183)

5x – 3y = 13 --(1)
4x – 3y = 11 --(2)
(1) + (2): 9x = 24
x = 24/9
Now, substitute x back into any one of the equations and isolate y to find it!